# Physics-informed neural networks with unknown measurement noise

### v4

@@ A.2 Noise forms
In Fig. 6, plots of the noise distributions defined in Section 5 are shown, together with examples of the PDFs learned by the EBM.
@@ A.3 Learning curves
In Fig. 7-8, learning curves for the experiments discussed in Sections 5.1-5.2, with Gaussian mixture (3G) noise, are provided. From the curves, it becomes apparent that PINN-EBM typically converges quickly to the correct solution, after the EBM has been initialized. In case of the exponential differential equation, PINN and PINN-off also converge fast, albeit towards often significantly less accurate solutions. For the Bessel function, both of them converge towards values close to the correct solution but do so only very slowly.
@@
Figure 8: Learning curves for the Navier Stokes equations (Section 5.2) with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
@@
A.4 Implementation details  [^3] The code for this project is available at https://github.com/ppilar/PINN-EBM.
The datasets employed in Sections 5.1 and A.1 contained 200 training points, 50 validation points and 2000 collocation points. For the Navier-Stokes example in Section 5.2, 4000 training points, 1000 validation points and 4000 collocation points were used. The collocation points were generated on a uniform grid. For both PINN and EBM, fully-connected neural networks with tanh activation function were used, with 4 layers of width 40 in case of the former, and 3 layers of width 5 for the latter. For the EBM, a dropout layer with factor $0.5$ was inserted before the last layer. Both inputs and outputs of the networks were normalized to their expected ranges. In case of the Navier-Stokes experiment, 5 layers of width 30 were used for the PINN.
The EBM was initialized after $i_{\rm EBM}=4000$ iterations, except for the Navier-Stokes example, where $i_{\rm EBM}=10000$ was chosen. The EBM was initialized for $N_{\rm EBM}=2000$ iterations. After the initialization, convergence was checked by confirming that the values of the PDF tended towards zero on both ends of the distribution; if the check failed, the initialization was repeated with an increased number of iterations until it passed.
The total number of iterations $N_{\rm tot}$ for the different experiments (Tables 1, 2, and 3; Fig. 7, 8, and 9) were $40.000$, $100.000$, and $100.000$, respectively, for the exponential, the Bessel, and the Navier Stokes equations. For the plots evaluating the impact of training parameters (Fig. 3, 10, 11, and 12), the corresponding values were $50.000$ for each of them. Unless stated otherwise, the weighting factor $\omega=1$ was chosen.
The Adam optimizer with learning rate of $0.002$ was used for both PINN and EBM. The batch size for data points was 200 (i.e. full batches were used for the exponential equation and the Bessel equation) and the batch size for collocation points was 100. Training the PINN-EBM typically took 30-50% longer than the standard PINN.
@@ Appendix B Impact of training parameters
In this appendix, we study the impact that various training parameters and dataset features have on the model performances.
@@ B.1 The number of training points
In Fig. 10, we investigate the impact of the number of training points available for the exponential differential equation (8) with Gaussian mixture noise. Starting with a small amount of training points $N_{d}=20$, all of the models perform similarly. As the number of training points increases, the model performances improve as well, although the improvement beyond $N_{d}=50$ seems negligible for PINN. The advantage of more training data is most pronounced for PINN-EBM. This makes sense as more training data will enable the EBM to learn a more accurate representation of the noise distribution. The results for smaller $N_{d}$ show that the EBM can learn reasonable noise distributions also for smaller data sets; thus, PINN-EBM does not depend on an outsized amount of training data to perform well.
Figure 10: Results for the exponential differential equation (8) with 3G noise as a function of the number of training points $N_{d}$. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
@@ B.2 The noise strength
Figure 11: Results for the exponential differential equation (8) with 3G noise as a function of the noise strength $f_{n}$. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
In Fig. 11, the impact of the noise strength on the results is investigated for the exponential differential equation (8) with Gaussian mixture noise. To vary the noise strength, the distributions depicted in Fig. 6 are rescaled by a factor $f_{n}$. It is apparent that PINN-EBM significantly outperforms PINN and PINN-off for high noise strengths. As the noise strength decreases, the performances of the different models converge. The NLL on the validation data remains a reliable indicator of which model performs best.
@@ B.3 The weighting factor $\omega$ for the PDE loss
In Fig. 12, results for different values of the weighting factor $\omega$ for the PDE loss in (6) and (7) are depicted. In case of PINN-EBM, the training starts with $\omega=1$ and the larger value of $\omega$ is only used as soon as the EBM has been initialized (compare Algorithm 4.3). The main questions that this investigation is supposed to answer are the following: do small values of the PDE residuals $f^{2}$ correspond to correct solutions? Is there an optimal value of $\omega$ and how can it be determined?
To answer the first question, it is apparent that $f^{2}$ becomes smaller as $\omega$ increases. However, the corresponding plots of the RMSE and $|\Delta\lambda|$ show that this does not imply correctness of the results. In case of the Bessel and the Navier Stokes equations, the results can become significantly worse as $\omega$ increases. This also shows that the PDE loss cannot, in general, override the detrimental effects of the non-Gaussian noise for the standard PINN. One may have assumed that the PDE would only be compatible with a very limited range of PDE parameters and hence for the PINN to be pushed towards the correct solution that way; however, this only seems to be the case for PINN-off when applied to the exponential differential equation.
Figure 12: The impact of the weighting factor $\omega$ on model performance is depicted for the different experiments considered in Section 5, with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
To answer the second question, when considering e.g. the plots for the RMSE, it is apparent that no such value $\omega$ exists. Only PINN seems to consistently perform better for small values $\omega$. To determine the best value of $\omega$ for a given case, the NLL on the validation data remains crucial. For PINN and PINN-off, worse predictions do not always translate into higher values of NLL, as the example of the Navier Stokes equations shows. In these cases, it is important to keep an eye on the learned values of $\lambda$ during training, as values $\lambda\approx 0$ can often indicate failure. For the PINN-EBM, however, there is a clear correlation between lower NLL and better results. Most importantly, the best solution as learned via PINN-EBM consistently results in lower NLL than PINN and PINN-off for any $\omega$.
ml>

### v5

@@ A.2 Noise Forms
In Fig. 6, plots of the noise distributions defined in Section 5 are shown, together with examples of the PDFs learned by the EBM. Overall, the EBM approximations are in good agreement with the true noise distributions.
@@ A.3 Learning Curves
In Figs. 7-9, learning curves for the experiments discussed in Sections 5.1-5.2 and Appendix A.1, with Gaussian mixture (3G) noise, are provided. From the curves, it becomes apparent that PINN-EBM typically converges quickly to the correct solution after the EBM has been initialized. In case of the exponential differential equation and the Navier-Stokes equations, PINN and PINN-off converge in a similar amount of iterations, albeit towards often significantly less accurate solutions. For the Bessel equation, both of them converge towards values close to the correct solution but do so only very slowly.
@@
Figure 8: Learning curves for the Navier-Stokes equations (Section 5.2) with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
@@
A.4 Additional Implementation Details  [^3] The code for this project is available at https://github.com/ppilar/PINN-EBM.
For the example of the Bessel equation A.1, the same parameters as for the exponential equation in Section 5.1 were used (see Section 5.3); only the total number of iterations for training the PINN was increased to $100.000$. For the plots evaluating the impact of training parameters (Figs. 10, 11, and 12), the total number of iterations was $50.000$ for all of the experiments. The noise distributions in Fig. 6 were scaled by factors of $f_{n}^{0}=1,0.03$, and $0.05$ for the exponential equation, the Bessel equation, and the Navier-Stokes equations, respectively.
To calculate the NLL for PINN and PINN-off, Gaussian distributions with the following parameters were employed: the mean was $\mu=0$ in case of PINN and $\mu=\theta_{0}$ in case of PINN-off. The standard deviations were estimated from the residuals between the training data and the model predictions.
@@ Appendix B Impact of Training Parameters
In this appendix, we study the impact that various training parameters and dataset features have on the model performances when considering the differential equations considered throughout the paper with 3G noise.
@@ B.1 The Number of Training Points
Fig. 10 depicts the impact of the number of data points available for training. The general trend is that the model performances improve with the number of training points. For PINN-EBM, the trend is most pronounced, which makes sense since more data points will allow it to learn a more accurate noise distribution. The main exception to this trend is PINN-off in case of the Bessel equation, where the volatility of the predictions increases with the number of training points. In terms of the accuracy of the PDE solution, as quantified via the RMSE and $|\Delta\lambda|$, the PINN-EBM consistently gives similarly good or better results than both PINN and PINN-off, independent of the number of training points. This shows that PINN-EBM does not depend on an outsized amount of training data to perform well.
Figure 10: The impact of the number of training points $N_{d}$ on the model performances is depicted for the different experiments considered in Section 5 and Appendix A.1, with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
@@ B.2 The Noise Strength
Figure 11: The impact of the noise strength $f_{n}$ on the model performances is depicted for the different experiments considered in Section 5 and Appendix A.1, with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
Figure 12: The impact of the weighting factor $\omega$ on the model performances is depicted for the different experiments considered in Section 5 and Appendix A.1, with 3G noise. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
In Fig. 11, the impact of the noise strength on the results is investigated. To vary the noise strength, the distributions depicted in Fig. 6 are rescaled by a factor of $f_{n}$ (after rescaling them with the factors $f_{n}^{0}$ corresponding to the different experiments, as given in Appendix A.4). It is apparent that PINN-EBM clearly outperforms PINN and PINN-off for high noise strengths. As the noise strength decreases, the performances of the different models converge: in case of the exponential equation and the Navier Stokes equations, the model performances are roughly the same in case of small noise. For the Bessel equation, however, the PINN-EBM retains a clear advantage regardless of the noise strength.
@@ B.3 The Weighting Factor $\omega$ for the PDE Loss
In Fig. 12, results for different values of the weighting factor $\omega$ for the PDE loss in (6) and (7) are depicted. In the case of PINN-EBM, the training starts with $\omega=1$ and the larger value of $\omega$ is only used as soon as the EBM has been initialized (compare Algorithm 4.3). This parameter deserves special attention since it can be tuned in real-world experiments and we are interested in the following question: is there an optimal value of $\omega$ and how can it be determined? To address this question, the PDE fulfillment $f^{2}$ and the NLL are of particular interest since only these quantities are available without knowing the correct PDE solution.
When considering $f^{2}$, it is apparent that it becomes smaller as $\omega$ increases. However, the corresponding plots of the RMSE and $|\Delta\lambda|$ show that this does not imply correctness of the results. In case of the Bessel and the Navier-Stokes equations, the results can become significantly worse as $\omega$ increases. This also shows that a more strongly weighted PDE loss cannot, in general, override the detrimental effects of the non-Gaussian noise for the standard PINN. One may have assumed that the PDE would only be compatible with a very limited range of PDE parameters and hence for the PINN to be pushed towards the correct solution that way; however, this only seems to be the case for PINN-off when applied to the exponential differential equation.
When considering e.g. the plots for the RMSE, it is apparent that no optimal value $\omega$ exists that would be valid for all the differential equations under consideration. Only PINN seems to consistently perform better for small values of $\omega$. To determine the best value of $\omega$ for a given case, the NLL on the validation data can be a useful criterion. For PINN and PINN-off, worse predictions do not always translate into higher values of NLL, as the example of the Navier-Stokes equations shows. In these cases, it is important to keep an eye on the learned values of $\lambda$ during training, as values $\lambda\approx 0$ can often indicate failure. For the PINN-EBM, however, there is a clear correlation between lower NLL and better results. Most importantly, when PINN-EBM achieves a better NLL than the other methods, the corresponding PDE solution is also better.